Fermi National Laboratory
 

Accelerator Update for December 8, 2006 - December 11, 2006

 

Friday December 8

The day shift began with Operations monitoring store 5113 (141.22E30), with the Antiproton Source (Pbar) stacking (58.6E10), with Linac and Booster operating normally, with Main Injector (MI) and Recycler in access, with NuMI and MiniBooNE off due to the MI access, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

A Booster expert began a study at 8:43 AM.

A power supply expert reported that he had replaced the regulator for a Pbar Debuncher Septa (D:ISEP).

Operations reported at 10:25 AM that the MI-10 upper power supply was back online.

MI came out of access around 11:11 AM.

The Neutron Therapy Facility (NTF) completed treating patients at 11:42 AM.

MiniBooNE resumed taking beam at 1:07 PM.

At 2:28 PM, Operations reported that 6 MI VCBs would not close. Experts investigated.

D:ISEP came into alarm at 2:36 PM. Operations called an expert who changed the Septa’s control card. It was back online by 3:52 PM.

NuMI resumed taking beam at 4:29 PM.

Operations tried to run beam to MI, but a Recycler kicker tripped off on a magnet temperature fault at 5:39 PM. It wouldn’t reset.

Kicker experts reported at 6:25 PM that the Recycler problem was in the MI/RR Tunnel. Operators prepared the area for access.

A kicker expert completed his repairs and was out of access by 7:46 PM.

At 7:55 PM, Operations reported that operators couldn’t close the MI VCBs. Operations called in experts.

Operations reestablished beam to MiniBooNE at 8:17 PM.

Experts closed the MI VCBs and Operations established beam to Pbar and NuMI at 9:38 PM.

The Recycler completed a stash of antiprotons at 10:49 PM. NuMI beam was again falling out of MI during the stash, but this stopped when Pbar resumed stacking.

Saturday December 9

The midnight shift began with Operations monitoring store 5113 (23.09E30) and a stash (65.9E10), with Pbar stacking (36.6E10), with Linac conducting studies, and with NuMI and MiniBooNE taking beam.

At 12:46 AM, a Linac RF station (LRF5) tripped off. A Linac expert was here and reported that the station was sparking. He thought that the switch tube was bad. Operations called in Linac technicians at 1:06 AM. The Linac expert found a loose BNC connector just as the techs arrived.

The Recycler completed its stash and Pbar resumed stacking at 2:27 AM.

Operations prepared the Recycler to stash and reported at 3:10 AM that the NuMI beam was falling out of MI during the setup.

At 3:18 AM, LRF2 tripped off. Luckily, the Linac techs were still here. They replaced a bad fuse for the switch tube filament at 3:49 AM.

At 3:58 AM, LRF2 tripped again. It reset okay.

At 4:46 AM, a Booster vacuum pump tripped off and caused two Booster RF stations to also trip (BRF7 & 8). Everything reset okay.

Operations terminated store 5113 at 6:53 AM.

At 7:04 AM, operators replaced a Tevatron (TeV) power supply in a sector F1 service building.

Operations began preparing the Recycler to stash at 7:03 AM. At 7:14 AM, Operations reported that half of the first transfer was lost and all of the second transfer due to a Recycler ACL script that had timed out. Operators managed to restart the script to complete the transfers.

At 7:34 AM, the MI beam permit tripped on an abort kicker ready indication. The permit reset, but NuMI beam began falling out of MI. An operator resolved this by making Time Line Generator (TLG) adjustments.

Operations put the TeV into shot setup at 7:53 AM.

At 8:04 AM, the Linac buncher tripped off. Operators had it back on about ten minutes later.

At 8:44 AM, the MI beam permit was tripping off due to the MI abort kickers. Operators went to MI-40 and adjusted the reservoir voltage per expert instruction. There was a shotgun like bang that shook the building. Operators called in experts. A capacitor had failed and damaged other equipment in a transformer cabinet.

Here are pictures of the remains of the capacitor and some of the resulting damage. Experts reported that they would have the repairs completed by 2 PM.

Operations reported at 9:42 AM that MiniBooNE and NuMI were taking beam and Pbar was stacking.

At 10:42 AM, Operations reported that a TeV F17 VCB wouldn’t close. Operations called in an expert.

At 11:09 AM, the TeV ramp dumped due to a F3 VCB opening. Operations called in another expert.

Operations made the CDF and D0 collision halls ready for an access at 11:18 AM.

CDF came out of access at 12:24 PM. D0 came out some time later.

Operations put the TeV through a QBS ramp at 1:52 PM.

Operations put the TeV into shot setup at 6:22 PM.

At 8:11 PM, Operations reported that the two last transfers of antiprotons were lost. This will be a 36x24 store (36 bunches of protons and 24 bunches of antiprotons).

A kicker expert turned on the MI-40 kickers at 8:31 PM.

Operations established store 5115 around 8:47 PM. D0 experienced difficulties reading their luminosity, so Operations extrapolated the initial luminosity to be around 78E30.

A Recycler expert reported at 9:25 PM that he thought the transfer loss was due to a MI low level RF problem. Operations called in a MI RF expert.

Operations reported at 10:20 PM that Pbar was stacking and NuMI was taking beam.

Sunday December 10

The midnight shift began with Operations monitoring store 5115 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

The Recycler completed a stash at 12:14 AM.

At 12:36 AM, Operations reported that LRF2 had lost its gradient for no apparent reason. This happened four more times over the next four hours. It reset each time.

The Recycler completed a stash at 3:05 AM.

At 4:56 AM, CDF reported having problems with their consoles, none of them were working. Operations contacted controls experts. CDF can’t take data without the ACNET network.

The Recycler completed a stash at 6:21 AM.

Operations reported at 8:08 AM that CDF had suffered a network switch failure.

At 8:33 AM, the Pelletron VME tripped off. An expert happened to be reading the electronic logbook from home, called, and suggested rebooting the electron cooling. The VME reset.

MI RF experts resumed their low-level investigations at 8:37 AM.

Operations reported at 9:24 AM that the H- Source had begun sparking near the end of the midnight shift.

Operations reported that CDF’s failed network switch was replaced at 9:56 AM.

A Preaccelerator expert instructed operators to adjust the H- cesium boiler voltage. He thought the temperature was too high.

The Recycler completed a stash at 11:19 AM.

At 11:40 AM, a MI octupole magnet (I:O124) tripped off on an over temperature indication and wouldn’t reset. This magnet normally runs at zero amps, so operators bypassed the alarm and contacted experts.

Operations terminated store 5115 at 12:32 PM and put the TeV into shot setup at 12:46 PM.

Operations established store 5116 at 2:26 PM with an initial luminosity of 122.10E30.

At 2:29 PM, Operations reported that the H- source started sparking again. An expert was called in.

The Recycler completed a stash at 3:15 PM.

At 4:10 PM, Operations reported a Pbar LCW leak alarm. Operations asked the Duty Electrician to investigate at CUB; the Duty Mechanic was busy.

At 4:52 PM, operators accessed the Pbar PreVault area (AP1) to look for a LCW leak because of its high humidity. There was no leak.

At 4:56 PM, a Preaccelerator expert accessed the H- Source to inspect cables.

Operations believed that the Pbar LCW leak was coming from the transport enclosure. Operations began preparing the Recycler to stash.

The Preaccelerator expert came out of access at 5:15 PM.

At 5:38 PM, a Booster East Anode Power Supply (EAPS) tripped off. Operators investigated and found that the supply’s yard breaker had tripped. An operator contacted an expert.

The Recycler completed a stash around 5:43 PM. Operations made Pbar ready for access.

Operations reported finding a broken LCW hose on a Pbar Debuncher quadrupole magnet (D:Q924) at 6:02 PM. They valved out the magnet and called in water experts. Power supply experts also had to replace the HS926 shunt

Operations reported that BRF2 had caused the EAPS yard breaker to trip. With the help of the Duty Electrician, Operations reset the breaker. BRF2 was disconnected.

Operations had established beam to MiniBooNE and NuMI by 7:42 PM.

At 7:52 PM, Operations reported that they could not reestablish stacking. Operators investigated.

Switchyard (SY) experts established beam to the SY absorber at 8:19 PM.

Operations reported at 8:34 PM that Pbar was stacking poorly. By 9:01 PM, operators had restored previous settings and resumed a decent stacking rate.

At 9:29 PM, a MI kicker (I:KPS4A) tripped off on a trigger ready indication. This was the same indication seen before the earlier failure this weekend. I:KPS4B tripped off because of 4A. They reset, but then tripped off again. At 10:56 PM, operations contacted an expert who had operators lower the kicker power supply peak voltage from 34.5kV to 34kV. He also suggested a few other changes, which the operators completed.

SY experts completed their commissioning for the night at 10:54 PM.

Monday December 11

The midnight shift began with Operations monitoring store 5116 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

The Recycler completed a stash at 12:45 AM.

At 2:13 AM, BRF19 tripped off. Operations reduced the beam intensity and bypassed the station.

The Recycler completed a stash at 4:11 AM.

The Recycler completed a stash at 7:22 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow cryo system experts to conduct wet engine maintenance at A3 and C2 after store 5116 gets terminated, which will be around noon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/28/2007    email Fermilab